OpenStudy (anonymous):

find dy/dx for (sec2x)/(x^2)

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Ive seen that before, but ti doesn't make sense. I need it in Lamens terms.

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

2sec2x(xtan2x−1)/x3

OpenStudy (anonymous):

I need to know how to solve it.

OpenStudy (anonymous):

first its rule the differentiation of f/g= (f’ g - g’ f )/g2

OpenStudy (anonymous):

ya, the product rule, right?

OpenStudy (anonymous):

then compare the rule wiyh ur problem f=sec2x g=x^2

OpenStudy (luigi0210):

Come on now, don't be afraid of the quotient rule.. it works better that way.

OpenStudy (anonymous):

no, I did use it, and I got the wrong answer.

OpenStudy (luigi0210):

Did you do it right is the question?

OpenStudy (anonymous):

I well, I did substituted into the rule right, but after that it gets kinda chaotic and i think i went wrong somewhere

OpenStudy (jdoe0001):

to get the derivative for sec(2x) you'd need to use the chain-rule I'd think

OpenStudy (jdoe0001):

other than that, is just a plain quotient rule

OpenStudy (luigi0210):

Here: \[\LARGE \frac{d}{dx}\frac {sec2x}{x^2}\] \[\LARGE =\frac{x^2*(2sec2xtan2x)-2x(sec2x)}{(x^2)^2}\]

OpenStudy (anonymous):

Let me try that quick, and I really appreciate your guys/girls time and help so far.

OpenStudy (luigi0210):

I know I might be wrong for \[\LARGE \frac{d}{dx} sec~2x\]Not sure..

OpenStudy (anonymous):

ummm, i have